# Dissipation meets conformal interface: How the relaxation rate is suppressed

### v1

@@ Abstract
Conformal interfaces play an important role in quantum critical systems. In closed systems, the transmission properties of conformal interfaces are typically characterized by two quantities: One is the effective central charge $c_{\text{eff}}$, which measures the amount of quantum entanglement through the interface, and the other is the transmission coefficient $c_{\text{LR}}$, which measures the energy transmission through the interface. In the present work, to characterize the transmission property of conformal interfaces in open quantum systems, we propose a third quantity $c_{\text{relax}}$, which is defined through the ratio of Liouvillian gaps with and without an interface. Physically, $c_{\text{relax}}$ measures the suppression of the relaxation rate towards a steady state when the system is subject to a local dissipation. We perform both analytical perturbation calculations and exact numerical calculations based on a free fermion chain at the critical point. It is found that $c_{\text{relax}}$ decreases monotonically with the strength of the interface. In particular, $0\leq c_{\text{relax}}\leq c_{\text{LR}}\leq c_{\text{eff}}$, where the equalities hold if and only if the interface is totally reflective or totally transmissive. Our result for $c_{\text{relax}}$ is universal in the sense that $c_{\text{relax}}$ is independent of (i) the dissipation strength in the weak dissipation regime and (ii) the location where the local dissipation is introduced. Comparing to the previously known $c_{\text{LR}}$ and $c_{\text{eff}}$ in a closed system, our $c_{\text{relax}}$ shows a distinct behavior as a function of the interface strength, suggesting its novelty to characterize conformal interfaces in open systems and offering insights into critical systems under dissipation.
@@ I Introduction
Impurities are commonly present in real physical systems and have a significant impact on both their equilibrium and non-equilibrium properties [1, 2, 3, 4, 5, 6, 7, 8]. In critical systems, when the impurity itself is a scale-invariant object, it not only encodes the universal properties of the bulk criticality, but also gives rise to fascinating phenomena that are invisible in the bulk [9, 10, 11, 12, 13, 14, 15, 16, 17]. Such impurities are known as conformal interfaces (or defects), which have attracted considerable attention due to their conceptual contribution to various fields, including the study of the Kondo effect [3, 9, 11], the development of renormalization group flows [18, 19, 20, 21, 22, 23, 24], the understanding of black hole evaporation [25, 26, 27], and applications to measurement problems [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45], among others.
For a closed quantum critical system in 1+1 dimensions, i.e., when the system is decoupled from environment, there are mainly two quantities to characterize a conformal interface: One is the effective central charge $c_{\text{eff}}$, and the other is the transmission coefficient $c_{\text{LR}}$. These two quantities characterize the amount of information transmission and the amount of energy transmission across the interface, respectively  [16, 17, 21, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. Both $c_{\text{eff}}$ and $c_{\text{LR}}$ are universal in characterizing the interface. For a general 1+1 dimensional conformal field theories (CFTs), no matter the system is in equilibrium or out of equilibrium, it is found that the quantum entanglement across the interface is described by the same $c_{\text{eff}}$ [54]. For $c_{\text{LR}}$, which characterizes the energy transmission, it was shown in [49] that $c_{\text{LR}}$ is independent of the details of initial state.
@@
Figure 1: Summary of the main results in this work. (a) A sketch for introducing a local dissipation to a one-dimensional quantum critical system in the presence of a conformal interface. The total system is of length $L=2N$, the conformal interface is located at $x=N$, and the dissipation is introduced at $x=n_{d}$. (b) The relaxation coefficient $c_{\text{relax}}$ and its comparison with $c_{\text{LR}}$ and $c_{\text{eff}}$, plotted as a function of the interface parameter $\lambda$ [see (7)] in a free-fermion lattice model at the critical point. The markers represent the exact numerical results of $c_{\text{relax}}$ for various choices of locations $n_{d}$ where the dissipation is introduced. For example, $n_{d}=1$ corresponds to the left boundary, and $n_{d}=N-1$ corresponds to the site next to the conformal interface. The green dashed line is the analytical result of $c_{\text{relax}}$ in Eq. (42), obtained from perturbative calculations. The blue and red lines are analytic results of $c_{\text{LR}}$ and $c_{\text{eff}}$ in Eqs. (12) and (13) respectively, with the corresponding lattice model calculations given in the appendix. Here the dissipation strength is $\gamma=0.05$, and the central charge of the bulk critical theory is $c=1$. Bottom left: A summary of defining properties for $c_{\text{eff}}$, $c_{\text{LR}}$, and $c_{\text{relax}}$.
@@ I.2 Main results
In addition, if the local dissipation is introduced at the boundary, we find that $c_{\text{relax}}$ in (4) holds even when the dissipation strength is out of the weak dissipation regime.
@@
The structure of the rest of this work is organized as follows. In Section II, we introduce the setup of adding local dissipation to a critical system with a conformal interface, along with a review of known results on energy and entanglement transmission in a closed quantum critical system. In Section III, we analyze the spectral structure of the Liouvillian using the third quantization formalism, by translating the problem to the study of an effective non-Hermitian Hamiltonian. Based on this, in Section IV, we perform a perturbative analysis of the Liouvillian gap, which is the main focus of this work. By combining analytical results with numerical calculations, we demonstrate a universal suppression of relaxation rate, with results presented in Section V. This allows us to define the relaxation coefficient $c_{\text{relax}}$ to characterize the conformal interface in a dissipative system. In Section VI, we conclude with a discussion and mention several future directions. There are also several appendices: We show how to extract $c_{\text{LR}}$ and $c_{\text{eff}}$ from the free fermion lattice models in Appendix A and Appendix B, respectively. Then in Appendix C, we give details on the perturbative calculation of the Liouvillian gap for arbitrary choices of the local dissipation location, as a supplement to the cases discussed in the main text. Finally, in Appendix D, we study the out-of-equilibrium quantum dynamics of the particle density, showing how the dissipative system approaches its steady state.
@@ II Setup and model
Figure 2: Two free-fermion chains of size $N$ are connected via a conformal interface characterized by a parameter $\lambda$. A local dissipation of strength $\gamma$ is introduced at site $n_{d}\in[1,N]$ on the left chain. The total system size is $L=2N$.
@@
[…]
@@ III Spectral structure of the Liouvillian: Third quantization formalism
[…]
@@
Figure 3: Rapidity spectrum $\{\beta_{i}\}$ for the free fermionic chain of length $L=2N=20$, with dissipation strength $\gamma=0.05$. (a) and (b): The local dissipation is added on the left boundary, i.e. $n_{d}=1$, and the interface parameter is set as (a) $\lambda=1$ (totally transmissive) and (b) $\lambda=0.4$ (partially transmissive). (c) and (d): The local dissipation is added right next to the interface, i.e. at $n_{d}=N-1$, and the interface parameter is set as (c) $\lambda=1$ (totally transmissive), and (d) $\lambda=0.4$ (partially transmissive). Orange dots represent rapidities with smallest real part other than zero, which determines the Liouvillian gap. Red dots represent rapidities with a vanishing real part and correspond to oscillating modes. Their appearance depends on the location of dissipation $n_{d}$ and the system size $2N$. All higher modes (decay modes) are denoted by blue dots.
[…]
@@
Figure 4: Full spectrum of the Liouvillian superoperator for a total size $L=2N=8$, with dissipation strength $\gamma=0.05$, and $\lambda=0.8$. The local dissipation is applied at (a) $n_{d}=1$ and (b) $n_{d}=N-1$. All eigenvalues of the Liouvillian come in complex conjugate pairs with non-positive real parts. Green dots show the steady-state solution. Red dots are the oscillatory modes with zero real part. Orange dots represent the slowest decaying modes, and blue dots represent all higher modes.
@@ IV The Liouvillian gap: from perturbation theory to numerics
[…]
@@
Figure 6: The Liouvillian gap $g(\lambda)$ as a function of the total system size $L=2N$, with local dissipation applied at $n_{d}=N-1$, the site right next to the interface. For various values of the interface parameter $\lambda$, we find an excellent agreement between the numerical results (dots) and the analytical results (dashed line) in Eq. (38). Different from the boundary dissipation case, the result here depends on whether $02N+1\bmod 3=$ 0. Here the total system size is chosen as $8002N=400,420,440,\dots,$ 800, and the dissipation strength is chosen as $\gamma=0.05$.
@@
Similar to the boundary dissipation case, for $n_{d}=N-1$, the Liouvillian gap also exhibits an asymptotic scaling behavior of $g\propto N^{-3}$, no matter there is a conformal interface or not. We compare this analytical solution with the numerical calculations in Fig. 6, and find an excellent agreement 444 As a remark, we want to point out that the case of $02N+1\bmod 3=$ 0 differs from the case of $2N+1\bmod 3\neq 0$ in hosting oscillatory coherence, i.e. time-dependent modes that never decay. Here, we are only interested in the relaxation rate of the decay modes. .
@@
V Suppression of relaxation for characterizing conformal interface

### v2

# Dissipation meets conformal interface in open quantum systems: How the relaxation rate is suppressed

@@ Abstract
Conformal interfaces play an important role in quantum critical systems. In closed systems, the transmission properties of conformal interfaces are typically characterized by two quantities: One is the effective central charge $c_{\text{eff}}$, which measures the amount of quantum entanglement through the interface, and the other is the transmission coefficient $c_{\text{LR}}$, which measures the energy transmission through the interface. In the present work, to characterize the transmission property of conformal interfaces in open quantum systems, we propose a third quantity $c_{\text{relax}}$, which is defined through the ratio of Liouvillian gaps with and without an interface. Physically, $c_{\text{relax}}$ measures the suppression of the relaxation rate towards a steady state when the system is subject to a local dissipation. We perform both analytical perturbation calculations and exact numerical calculations based on a free fermion chain at the critical point. It is found that $c_{\text{relax}}$ decreases monotonically with the strength of the interface. In particular, $0\leq c_{\text{relax}}\leq c_{\text{LR}}\leq c_{\text{eff}}$, where the equalities hold if and only if the interface is totally reflective or totally transmissive. Our result for $c_{\text{relax}}$ is universal in the sense that $c_{\text{relax}}$ is independent of (i) the dissipation strength in the weak dissipation regime and (ii) the location where the local dissipation is introduced. When local dissipation is introduced at the boundary, we analytically show that the result holds for any finite dissipation strength; when dissipation is introduced in the bulk or near the interface, we perform a perturbative analysis to establish the relation’s validity in generic settings. Comparing to the previously known $c_{\text{LR}}$ and $c_{\text{eff}}$ in a closed system, our $c_{\text{relax}}$ shows a distinct behavior as a function of the interface strength, suggesting its novelty to characterize conformal interfaces in open systems and offering insights into critical systems under dissipation.
@@ I Introduction
Impurities are commonly present in real physical systems and have a significant impact on both their equilibrium and non-equilibrium properties  [1, 2, 3, 4, 5, 6, 7, 8]. In critical systems, when the impurity itself is a scale-invariant object (a conformal defect/interface), it not only encodes the universal properties of the bulk criticality, but also gives rise to fascinating phenomena that are invisible in the bulk [9, 10, 11, 12, 13, 14, 15, 16, 17]. Such impurities have attracted considerable attention due to their conceptual contribution to various fields, including the study of the Kondo effect [3, 9, 11], the development of renormalization group flows [18, 19, 20, 21, 22, 23, 24], the understanding of black hole evaporation [25, 26, 27], and applications to measurement problems [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45], among others.
Because a conformal interface introduces no intrinsic length scale, the reflection and transmission coefficients are independent of the magnitude of the incident momentum or energy. This independence is a useful microscopic diagnostic of scale invariance of a conformal interface and, in closed quantum critical systems, provides a characterization via the transmission coefficient $c_{\text{LR}}$ [16, 17, 21, 46, 47, 48], which quantifies energy transmitted across the interface. Another widely used characterization of conformal interfaces in closed system (mainly 1+1-dimensional) is the effective central charge $c_{\text{eff}}$ extracted from entanglement entropy, which measures the amount of information transmitted across the interface [49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. Both $c_{\text{eff}}$ and $c_{\text{LR}}$ are universal in characterizing the interface. For a general 1+1 dimensional conformal field theories (CFTs), no matter the system is in equilibrium or out of equilibrium, it is found that the quantum entanglement across the interface is described by the same $c_{\text{eff}}$ [54]. For $c_{\text{LR}}$, which characterizes the energy transmission, it was shown in [46] that $c_{\text{LR}}$ is independent of the details of initial state.
@@
Figure 1: Summary of the main results in this work. (a) A sketch for introducing a local dissipation to a one-dimensional quantum critical system in the presence of a conformal interface. The total system is of length $L=2N$, the conformal interface is located at $x=N$, and the dissipation is introduced at $x=n_{d}$. (b) The relaxation coefficient $c_{\text{relax}}$ and its comparison with $c_{\text{LR}}$ and $c_{\text{eff}}$, plotted as a function of the interface parameter $\lambda$ [see (7)] in a free-fermion lattice model at the critical point. The markers represent the exact numerical results of $c_{\text{relax}}$ for various choices of locations $n_{d}$ where the dissipation is introduced (Note: markers appear on the top of each other in the figure). For example, $n_{d}=1$ corresponds to the left boundary, and $n_{d}=N-1$ corresponds to the site next to the conformal interface. The green dashed line is the analytical result of $c_{\text{relax}}$ in Eq. (56), obtained from perturbative calculations. The blue and red lines are analytic results of $c_{\text{LR}}$ and $c_{\text{eff}}$ in Eqs. (12) and (13) respectively, with the corresponding lattice model calculations given in the appendix. Here the dissipation strength is $\gamma=0.05$, and the central charge of the bulk critical theory is $c=1$. Bottom left: A summary of defining properties for $c_{\text{eff}}$, $c_{\text{LR}}$, and $c_{\text{relax}}$.
@@ I.2 Main results
In addition, if the local dissipation is introduced at the boundary, we find that $c_{\text{relax}}$ in (4) holds even when the dissipation strength is out of the weak dissipation regime (see section V).
@@
The structure of the rest of this work is organized as follows. In Section II, we introduce the setup of adding local dissipation to a critical system with a conformal interface, along with a review of known results on energy and entanglement transmission in a closed quantum critical system. In Section III, we analyze the spectral structure of the Liouvillian using the third quantization formalism, by translating the problem to the study of an effective non-Hermitian Hamiltonian. Based on this, in Section IV, we perform a perturbative analysis of the Liouvillian gap, which is the main focus of this work. In Sec. V, we analytically compute the Liouvillian gap for finite dissipation strength using a plane-wave ansatz, assuming that local dissipation is introduced at the boundary. By combining analytical results with numerical calculations, we demonstrate a universal suppression of relaxation rate, with results presented in Section VI. This allows us to define the relaxation coefficient $c_{\text{relax}}$ to characterize the conformal interface in a dissipative system. In Section VII, we conclude with a discussion and mention several future directions. There are also several appendices: We show how to extract $c_{\text{LR}}$ and $c_{\text{eff}}$ from the free fermion lattice models in Appendix A and Appendix B, respectively. Then in Appendix C, we give details on the perturbative calculation of the Liouvillian gap for arbitrary choices of the local dissipation location, as a supplement to the cases discussed in the main text. Finally, in Appendix D, we study the out-of-equilibrium quantum dynamics of the particle density, showing how the dissipative system approaches its steady state.
@@ II Setup and model
Figure 2: Two free-fermion chains of size NN are connected via a conformal interface characterized by a parameter $\lambda$. A local dissipation of strength $\gamma$ is introduced at site $n_{d}\in[1,N]$ on the left chain. The total system size is $L=2N$.
@@
where $\rho(t)$ is the density matrix of the system at time $t$, and $\mathcal$$\widehat${L}} is called the Liouvillian (also called Lindbladian) superoperator that generates the dissipative dynamics. Here $H$ is the Hamiltonian in Eq.(6) and(7) that drives the unitary dynamics. For the jump operators $L_{\mu}$, we consider the following choice :
@@ III Spectral structure of the Liouvillian: Third quantization formalism
In the third quantization formalism, the Liouvillian $\mathcal$$\widehat${L}} is treated as a linear map over a space of operators of dimension $4^{2N}$, where each vector corresponds to an operator of $4N$ Majorana fermions. The inner product of this operator space is defined as $\langle A|B\rangle=2^{-2N}\operatorname{Tr}(A^{\dagger}B)$, where $2^{-2N}$ is the normalization factor. In this operator space, the Liouvillian can be re-expressed in terms of $8N$ Hermitian Majorana fermionic maps $\widehat{a}_{i}$ as  [^2] It is reminded that here we have $8N$ maps instead of $4N$, because the original complex-fermion chain is of length $L=2N$ instead of $N$.
@@
Figure 3: Rapidity spectrum $\{\beta_{i}\}$ for the free fermionic chain of length $L=2N=20$, with dissipation strength $\gamma=0.05$. (a) and (b): The local dissipation is added on the left boundary, i.e. $n_{d}=1$, and the interface parameter is set as (a) $\lambda=1$ (totally transmissive) and (b) $\lambda=0.4$ (partially transmissive). (c) and (d): The local dissipation is added right next to the interface, i.e. at $n_{d}=N-1$, and the interface parameter is set as (c) $\lambda=1$ (totally transmissive), and (d) $\lambda=0.4$ (partially transmissive). Orange dots represent rapidities with smallest real part other than zero, which determines the Liouvillian gap. Green dots represent rapidities with a vanishing real part and correspond to oscillating modes. Their appearance depends on the location of dissipation $n_{d}$ and the system size $2N$. All higher modes (decay modes) are denoted by blue dots.
Before applying the above procedure to our setup, let us briefly introduce several key properties and concepts about the Liouvillian spectrum $\{\Lambda_{i}\}$. First, all $\Lambda_{i}$ satisfy $\operatorname{Re}(\Lambda_{i})\leq 0$. Second, a non-equilibrium steady state is an eigenstate of the Liouvillian $\mathcal$$\widehat${L}} with eigenvalue 0. This steady state is time-independent and free from decoherence. Third, in certian cases, there are modes with nonzero purely imaginary eigenvalues, which is a phenomenon known as oscillatory coherence [71, 77, 78]. These modes are time-dependent but free from decay. In this case, depending on the initial state, the system does not necessarily relax to a steady state. This is totally fine, since in this work we are mainly interested in the relaxation time, i.e, the longest lifetime of the decay modes (all the modes other than the steady state and the oscillating modes).
@@
Figure 4: Full spectrum of the Liouvillian superoperator for a total size $L=2N=8$, with dissipation strength $\gamma=0.05$, and $\lambda=0.8$. The local dissipation is applied at (a) $n_{d}=1$ and (b) $n_{d}=N-1$. All eigenvalues of the Liouvillian come in complex conjugate pairs with non-positive real parts. Purple dots show the steady-state solution. Green dots are the oscillatory modes with zero real part. Orange dots represent the slowest decaying modes, and blue dots represent all higher modes.
@@ IV The Liouvillian gap: from perturbation theory to numerics
Figure 5: The Liouvillian gap $g(\lambda)$ as a function of the total system size $L=2N$, with local dissipation applied at the left boundary ($n_{d}=1$). For various values of the interface parameter $\lambda$, there is an excellent agreement between the numerical results (dots) and the analytical result from perturbative analysis (dashed line) in Eq. (35). Here the total system size is chosen as $L=2N=200,\,260,\,320,\dots,800$, and the dissipation strength is set to be $\gamma=0.05$.
@@
Figure 6: The Liouvillian gap $g(\lambda)$ as a function of the total system size $L=2N$, with local dissipation applied at $n_{d}=N-1$, the site right next to the interface. For various values of the interface parameter $\lambda$, we find an excellent agreement between the numerical results (dots) and the analytical results (dashed line) in Eq. (38). Different from the boundary dissipation case, the result here depends on whether $2N+1\bmod 3=0$. Here the total system size is chosen as $2N=400,420,440,\dots,800$, and the dissipation strength is chosen as $\gamma=0.05$.
@@
Similar to the boundary dissipation case, for $n_{d}=N-1$, the Liouvillian gap also exhibits an asymptotic scaling behavior of $g\propto N^{-3}$, no matter there is a conformal interface or not. We compare this analytical solution with the numerical calculations in Fig. 6, and find an excellent agreement  [^4] As a remark, we want to point out that the case of $2N+1\bmod 3=0$ differs from the case of $2N+1\bmod 3\neq 0$ in hosting oscillatory coherence, i.e. time-dependent modes that never decay. Here, we are only interested in the relaxation rate of the decay modes. .
@@
V Liouvillian gap for boundary dissipation with finite dissipation strength
In the previous section, working in the perturbative regime of the dissipation strength $\gamma$, we derived the relation between the Liouvillian gap and the interface parameter $\lambda$ for an arbitrary location of the additional local dissipation. Here, we show that when the dissipation is introduced exactly at the boundary, the asymptotic behavior of the Liouvillian gap at large system sizes can be obtained at finite $\gamma$, rendering the conclusion non-perturbative.
Our derivation follows the approach in Ref. [64, 65]. We consider the eigenvalue problem $\textbf{H}_{\text{eff}}\bm{\psi}=E\bm{\psi}$ which satisfy following conditions. The effective Hamiltonian $\textbf{H}_{\text{eff}}$ has the same expression as (31). (i) In the bulk, the eigenvalue equation is given by
$J\psi_{i-1}+J\psi_{i+1}=E\psi_{i}.\\$ (40)
(ii) At the two boundaries, the eigenvalue problem satisfies
$\left\{\begin{split}-i\dfrac{\gamma}{2}\psi_{1}+J\psi_{2}=&E\psi_{1}\\ J\psi_{2N-1}=&E\psi_{2N}\end{split}\right$. (41)
Thus, one can introduce plane wave ansatz similar to [64]
$\psi_{j}=Az^{j}+Bz^{-j},$ (42)
where $z\in\mathbb{C}.$, and the eigenvalue can be obtained as
$E=-\frac{1}{2}(z+z^{-1}).$ (43)
However, in this work we also have a defect at the middle of the system thus it modifies the quantization. (iii) Eigenvalue problem should also satisfy additional conditions due to the defect at the middle of the chain
$-\frac{1}{2}\psi_{N-1}+\dfrac{\sqrt{1-\lambda^{2}}}{2}\psi_{N}-\dfrac{\lambda}{2}\psi_{N+1}=E\psi_{N},$ (44)
$-\dfrac{\lambda}{2}\psi_{N}-\dfrac{\sqrt{1-\lambda^{2}}}{2}\psi_{N+1}-\frac{1}{2}\psi_{N+2}=E\psi_{N+1}.$ (45)
So, in this work we introduce the following plane wave ansatz
$\psi_{j}(k)=\begin{cases}\alpha_{k}(Az^{j}+Bz^{-j})\qquad\quad 1\leqslant j\leqslant N,\\ \beta_{k}(Cz^{j}+Dz^{-j})\quad N+1\leqslant j\leqslant 2N,\end{cases}$ (46)
[…]
$\begin{split}&A(1-i\gamma z)+B(1-i\gamma z^{-1})=0,\\ &A(z^{N+1}\alpha_{k}+z^{N}\alpha_{k}\delta)+B(z^{-N-1}\alpha_{k}+z^{-N}\alpha_{k}\delta)+\\ &C(-z^{N+1}\beta_{k}\lambda)+D(-z^{-N-1}\beta_{k}\lambda)=0,\\ &A(-z^{N}\alpha_{k}\lambda)+B(-z^{-N}\alpha_{k}\lambda)+C(z^{N}\beta_{k}-z^{N+1}\beta_{k}\delta)+\\ &D(z^{-N}\beta_{k}-z^{-L-1}\beta_{k}\delta)=0,\\ &C(z^{2N+1})+D(z^{-2N-1})=0.\end{split}$ (47)
where $\delta=\sqrt{1-\lambda^{2}}$. The above equations have non-trivial solutions when $\theta_{k}$ satisfy the following quantization condition
$\text{sin}(2N+1)\theta_{k}-i\gamma\text{sin}(2N)\theta_{k}+i\gamma\sqrt{1-\lambda^{2}}\text{sin}\theta_{k}=0.$ (48)
We know that for the boundary dissipation Liouvillian gap corresponds to $k=1$ case. Now, similar to previous works [64, 65], we set $\theta_{1}=\dfrac{\pi}{(2N+1)}+\dfrac{\theta_{R}+i\theta_{I}}{(2N+1)^{2}}$. After substituting $\theta_{1}$ in (48) and keeping terms upto leading order terms in $1/N$, we get
$\theta_{R}=\dfrac{\pi\gamma^{2}(1-\sqrt{1-\lambda^{2}})}{1+\gamma^{2}},$ (49)
$\theta_{I}=-\dfrac{\pi\gamma(1-\sqrt{1-\lambda^{2}})}{1+\gamma^{2}}.$ (50)
Now, we can calculate the eigenvalue
$\begin{split}E=&-\frac{1}{2}(z+z^{-1})\\ =&-\dfrac{1}{2}\text{exp}\left(\dfrac{i\pi}{2N+1}\right)\left(1-\dfrac{\theta_{I}}{2N+1}\right)\\ &-\dfrac{1}{2}\text{exp}\left(\dfrac{-i\pi}{2N+1}\right)\left(1+\dfrac{\theta_{I}}{2N+1}\right)\\ =&-\text{cos}\left(\dfrac{\pi}{2N+1}\right)+\dfrac{i\theta_{I}}{(2N+1)}\text{sin}\left(\dfrac{\pi}{2N+1}\right).\end{split}$ (51)
Substituting $\theta_{I}$ from (50), we get the maximum imaginary part of eigenvalues as [^6] For the Liouvillian gap $g=-2\max_{\operatorname{Im}[E_{i}]\neq 0}\left\{\operatorname{Im}[E_{i}]\right\}$, where $E_{i}$ are the eigenvalues of Z. The effective Hamiltonian matrix is related to the matrix Z as $\textbf{Z}=\frac{1}{2}\textbf{H}_{\text{eff}}$ (See (31)). Hence, the Liouvillian gap is $g=-\text{max}\{E_{\text{imag}}\}$.
$\text{max}\{E_{\text{imag}}\}=-\dfrac{\pi\gamma(1-\sqrt{1-\lambda^{2}})}{(1+\gamma^{2})(2N+1)^{2}}\text{sin}\left(\dfrac{\pi}{2N+1}\right).$ (52)
Thus, for $2N>>1$, we find the expression of Liouvillian gap for finite values of dissipation strength $\gamma$ as
$g(\lambda)=-\text{max}\{E_{\text{imag}}\}=\dfrac{\gamma\pi^{2}(1-\sqrt{1-\lambda^{2}})}{(1+\gamma^{2})(2N+1)^{3}}.$ (53)
Finally, we note that for local dissipation away from the boundary, the plane-wave ansatz in Eq. (46) must be modified, which in turn alters the quantization condition in Eq. (48). In this case, the Liouvillian gap does not necessarily correspond to the $k=1$ mode.
VI Suppression of relaxation for characterizing conformal interface